Circular No. 6097

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA CASSIOPEIAE 1993 S. J. Boyle, University of London Observatory, writes: “Images of N

Cas 1993 were obtained by M. Pontefract and K. Siebert with the 0.6-m Radcliffe refractor (+ CCD) at the University of London Observatory on Oct. 4. A slightly asymmetric region of emission was detected, with a radius of ∼ 9′′. This region is larger than predicted by any plausible rate of expansion of the shell produced by the nova’s outburst on 1993 Dec. 4. We attribute the emission to the excitement of material previously ejected from the progenitor star and estimate that the episode of mass-loss occurred at some epoch between 1920 and 1960. Spectra in the range 540–700 nm obtained simultaneously by J. Bly and P. Newman with the 0.6-m Allen reflector show emission lines due to Hα, O I (630.0 nm), and N II (575.5 nm). S. Boyle measured the HWZI of the last-noted line to be 970 km/s.”

POSSIBLE GRAVITATIONAL MICROLENSING EVENT E. Giraud, European Southern Observatory, writes: “I have obtained

two exposures, with the New Technology Telescope at La Silla on the night of Oct. 18–19, of the microlens candidate in the Large Magellanic Cloud reported by the MACHO collaboration (IAUC 6095). Images were obtained with a Bessel R filter (central wavelength 645.0 nm; FWHM 155.0 nm). I also obtained a spectrum (range 600–800 nm) on the same night. The microlens candidate has a faint companion to the east; this explains why the reference image provided in the anonymous FTP account (ibid.) appears to be elongated. The difference in magnitude of the lens candidate between the reference frame and the present observation is 0.89 ± 0.2 mag; the accuracy in relative photometry of the present measurement is 0.06 mag, while that of the reference frame is of the order of 0.2 mag. (A better determination of the zero point, coadding reference frames, can considerably reduce the error.) The spectrum shows a weak detection of Hα in absorption.”

PERIODIC COMET BORRELLY (1994l) Total visual magnitude estimates by C. S. Morris (southern California,

0.26-m reflector): Oct. 2.47 UT, 9.0; 3.47, 8.9; 7.47, 8.8; 8.48, 8.8; 9.48, 8.8; 12.44, 9.0.

1994 October 20 Daniel W. E. Green